Please write an equation that satisfies these conditions:
Parallel to 2x-y=6 and passes through (-6,-9)
Please show step by step!!
Thank you :)

Answer by ewatrrr(24785)


Hi
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
Line Parallel Parallel to 2x-y=6 and passes through (-6,-9)
2x-y=6 OR y = 2x-6 ( m = 2)
NEw Line: parallel Lines have identical slopes
y = 2x + b |Using ordered pair (-6,-9) to solve for b
-9 = 2*-6 + b
3 = b
y = 2x + 3